Sam reaches into a bag that contains 6 tiles numbered 1 through 6 and draws 2 stimultaneously. What is the probability that Sam drew 2 odd numbers?
Answer by stanbon(75887)

Sam reaches into a bag that contains 6 tiles numbered 1 through 6 and draws 2 stimultaneously. What is the probability that Sam drew 2 odd numbers?
--------------
There are three odd numbers from 1 to 6; you want 2; 3C2 = 3
There are 6C2= 15 different 2-tile results
P(2 odd)= 3/15 = 1/5 = 0.2
==============================
Cheers,
Stan H.